# code[Code][List of Codes]

### v1

@@ Abstract
Fine-grained and contact-rich manipulation remain challenging for robots, largely due to the underutilization of tactile feedback. To address this, we introduce TouchGuide, a novel cross-policy visuo-tactile fusion paradigm that fuses modalities within a low-dimensional action space. Specifically, TouchGuide operates in two stages to guide a pre-trained diffusion or flow-matching visuomotor policy at inference time. First, the policy produces a coarse, visually-plausible action using only visual inputs during early sampling. Second, a task-specific Contact Physical Model (CPM) provides tactile guidance to steer and refine the action, ensuring it aligns with realistic physical contact conditions. Trained through contrastive learning on limited expert demonstrations, the CPM provides a tactile-informed feasibility score to steer the sampling process toward refined actions that satisfy physical contact constraints. Furthermore, to facilitate TouchGuide training with high-quality and cost-effective data, we introduce TacUMI, a data collection system. TacUMI achieves a favorable trade-off between precision and affordability; by leveraging rigid fingertips, it obtains direct tactile feedback, thereby enabling the collection of reliable tactile data. Extensive experiments on five challenging contact-rich tasks, such as shoe lacing and chip handover, show that TouchGuide consistently and significantly outperforms state-of-the-art visuo-tactile policies.
@@ I Introduction
To address this fundamental limitation, we introduce TouchGuide, a novel cross-policy visuo-tactile fusion paradigm. Departing from conventional computer vision approaches that focus on “fusing multi-modal data per se”, we rethink the problem from a robotic control perspective. Our core idea is to extract task-relevant semantics from each modality and perform fusion within the low-dimensional action space in a way that can benefits robotic manipulation. Concretely, TouchGuide steers a pre-trained visuomotor policy based on diffusion or flow matching (e.g., the action expert in $\pi_{0.5}$) during inference. The process begins by generating a visually feasible, coarse action from visual observation alone in the early sampling phase. Then, in the later phase, a task-specific Contact Physical Model (CPM) injects tactile guidance, refining the action to better satisfy real-world physical contact constraints. The CPM is trained via contrastive learning on limited expert demonstrations to directly output a feasibility score. This score assesses whether the action comply with the contact physical constraints implied by the current tactile observations in task-specific environment, thereby generating a more robust steering signal. Notably, CPM can be applied across multiple policies, including VLAs [32, 5, 50, 58], diffusion policies [14, 69, 56, 44], and flow-matching policies [34, 71].
Moreover, in fine-grained and contact-rich manipulation, policy execution performance depends heavily on the quality of expert demonstrations. Teleoperation, due to high latency and indirect haptic feedback, hinders intuitive control. While handheld devices such as UMI [15] provide a platform-decoupled solution for data collection, existing systems relying on SLAM-based localization or motion capture (MoCap) cannot simultaneously achieve both high accuracy and low cost. To this end, we introduce TacUMI, designed according to the following principles: (1) Cost-effective (\$720 without tactile sensors) and lightweight (540g), while providing the high-precision data required for fine-grained manipulation. (2) Direct tactile feedback enabled by rigid fingertips. (3) A unified interface for synchronized acquisition of high-resolution visual and tactile signals.
@@ II Related Work
Figure 2: Overview of TacUMI data collection system. TacUMI (Collection-side) uses a Vive tracker for localization to obtain accurate end-effector poses, while the operator receives direct tactile feedback (Left). During policy inference, we use an execution-side device that is structurally identical to the collection-side TacUMI, coupled to different robot arms via an adapter.
@@ III-A System Overview
Accuracy and Low-cost. TacUMI is localized using a Vive tracker with two Lighthouse base stations, providing reliable and accurate pose estimates. In contrast, UMI [15] relies on SLAM-based tracking; for fine-grained tactile manipulation tasks which requires high precision, its accuracy is insufficient and tracking can be easily lost in feature-sparse scenes. Notably, a setup with one Vive tracker and two Lighthouse base stations costs only about \$720 (without tactile sensors), maintaining a low cost while providing accurate localization. A detailed comparison of data collection quality across different methods will be presented in the experiments (See Sec. V-D).
@@
Usability and Synchronization. The entire gripper adopts a unified hardware-software interface. By connecting it to a computer via a single aviation cable, we can stream observations, including tactile and visual signals as well as proprioceptive states. In addition, our data collection is compatible with the LeRobot [7] format and uses a unified software interface to acquire all modalities, thereby achieving effective software-level synchronization. Moreover, TacUMI captures high-resolution visual $\mathbf{V}_{t}\in\mathbb{R}^{640\times 480\times 3}$ and tactile $\mathbf{T}_{t}\in\mathbb{R}^{200\times 350\times 3}$ images at 30Hz, which suffices for the vast majority of fine-grained manipulation tasks. Meanwhile, we integrate an RK3576 CPU inside TacUMI, which enables on-board computation of force while capturing tactile images. Due to the modality-specific nature of tactile sensing (i.e., different tactile sensors cannot share the same tactile images for training), this capability is crucial for tactile data collection.
More details on hardware selection, design choices, and system parameters are provided in the Appendix -D.
@@ III-B Data Processing for Training
$\mathbf{p}_{\text{rel}}^{(i)}$ $={\mathbf{R}_{\text{ee}}^{(0)\top}}(\mathbf{p}_{\text{ee}}^{(i)}-\mathbf{p}_{\text{ee}}^{(0)})$ (1)
$\mathbf{R}_{\text{rel}}^{(i)}$ $={\mathbf{R}_{\text{ee}}^{(0)\top}}\mathbf{R}_{\text{ee}}^{(i)}$ (2)
@@ IV Method: TouchGuide
In this section, we introduce TouchGuide, a two-stage cross-policy multimodal fusion paradigm based on classifier guidance [18] (Sec. IV-A), inspired by recent advances in inference-time steering [51, 22, 55]. TouchGuide performs inference-time steering on a pre-trained diffusion-based or flow-matching-based (e.g., action expert of $\pi_{0.5}$) visuomotor policy $\pi_{\theta}(\mathbf{A}_{t}\,|\,\mathbf{V}_{t})$. During the early phase of the denoising or flow-matching process, the policy generates an initial coarse action based solely on visual observation, ensuring visual feasibility without yet enforcing tactile constraints. In the later phase, a task-specific Contact Physical Model (CPM, detailed in Sec. IV-B) provides tactile guidance via a feasibility score, which steers the action toward better satisfaction of real-world physical contact. The complete inference-time guidance process of TouchGuide is detailed in Sec. IV-C.
@@ IV-A Preliminaries: Classifier Guidance
$\hat{\epsilon}_{\theta}(x_{t})=\epsilon_{\theta}(x_{t})-\eta\sqrt{1-\bar{\alpha}_{t}}\nabla_{x_{t}}\log p_{\phi}(y\mid x_{t})$ (3)
@@
$\hat{u}_{\theta}(x_{t})=u_{\theta}(x_{t})+\eta\frac{t}{1-t}\nabla_{x_{t}}\log p_{\phi}(y\mid x_{t})$ (4)
@@ IV-B Contact Physical Model
18:     $\hat{u}\leftarrow u+\eta\frac{k}{1-k}\nabla_{\mathbf{A}_{t}^{k}}\mathbf{s}$
@@
At time step $t$, Contact Physical Model (CPM, shown in Fig. 3(a)) outputs a task-specific feasibility score conditioned on the tactile feedback $\mathbf{T}_{t}$, the visual observation $\mathbf{V}_{t}$ and the coarse action $\mathbf{A}_{t}$ generated by the early phase of pre-trained policy $\pi_{\theta}(\mathbf{A}_{t}\,|\,\mathbf{V}_{t})$. Specifically, we use pretrained DINOv2 as both the tactile and visual encoder to obtain the tactile and visual embedings, i.e., $\mathbf{T}_{t}^{\text{emb}}$ and $\mathbf{V}^{\text{emb}}_{t}$ respectively, as it provides effective latent-space representations of robotic observations [22, 78]. Then, we fuse these two embedings using N-layer Transformer encoder $\mathtt{TranEnc}(\cdot)_{\times\text{N}}$ and finally, yield the latent observation $\mathbf{O}_{t}$:
$\mathbf{O}_{t}^{\text{emb}}=\mathtt{TranEnc}(\mathtt{concat}[\mathbf{T}_{t}^{\text{emb}},\mathbf{V}_{t}^{\text{emb}}])_{\times\text{N}}\xrightarrow[]{\text{L2 Norm}}\mathbf{O}_{t}$ (5)
Similarly, we can obtain the latent action $\mathbf{a}_{t}$ by:
$\mathbf{A}_{t}^{\text{emb}}\xrightarrow[]{\text{L2 Norm}}\mathbf{a}_{t}$ (6)
where, $\mathbf{A}_{t}^{\text{emb}}$ is the action embedding generated by an action encoder taking $\mathbf{A}_{t}$ as input. The feasibility score $\mathbf{s}$ represents the cosine similarity between the observation and the action in the latent space [11], and is defined as follows:
$\mathbf{s}=s_{\phi}(\mathbf{V}_{t},\mathbf{T}_{t},\mathbf{A}_{t})=\mathbf{O}_{t}^{\top}\mathbf{a}_{t}$ (7)
$\mathcal{L}_{\text{CPM}}(\mathbf{O}_{t},\mathbf{a}_{t})=\frac{1}{2}(\mathcal{L}_{\mathbf{O}\to\mathbf{a}}+\mathcal{L}_{\mathbf{a}\to\mathbf{O}})$ (8)
$\mathcal{L}_{\mathbf{O}\to\mathbf{a}}$ $=-\frac{1}{M}\sum_{i=1}^{M}\log\frac{\exp(\mathbf{O}_{i}^{\top}\mathbf{a}_{i}/\tau)}{\sum_{j=1}^{M}\exp(\mathbf{O}_{i}^{\top}\mathbf{a}_{j}/\tau)}$ (9)
$\mathcal{L}_{\mathbf{a}\to\mathbf{O}}$ $=-\frac{1}{M}\sum_{i=1}^{M}\log\frac{\exp(\mathbf{a}_{i}^{\top}\mathbf{O}_{i}/\tau)}{\sum_{j=1}^{M}\exp(\mathbf{a}_{i}^{\top}\mathbf{O}_{j}/\tau)}$ (10)
where $\tau$ is a learnable temperature parameter and $M$ denotes the number of samples used during training. In order to enable the CPM to effectively operate in a noisy action space (generated by the early phase of the pretrained policy) during inference, we enhance the ground-truth action $\mathbf{A}_{t}$ by adding noise following a geometric distribution using the same noise scheduler as DP [22]; for flow-matching policies, we instead apply linear interpolation.
@@ IV-C Steering the Policies via Touch Guidance
$\hat{\epsilon}_{\theta}(\mathbf{A}_{t}^{k},\mathbf{O}_{t})=\epsilon_{\theta}(\mathbf{A}_{t}^{k},\mathbf{V}_{t})-\eta\sqrt{1-\bar{\alpha}_{k}}\nabla_{\mathbf{A}_{t}^{k}}s_{\phi}(\mathbf{O}_{t},\mathbf{A}_{t}^{k})$ (11)
$\hat{u}_{\theta}(\mathbf{A}_{t}^{k},\mathbf{O}_{t})=u_{\theta}(\mathbf{A}_{t}^{k},\mathbf{V}_{t})+\eta\frac{k}{k-1}\nabla_{\mathbf{A}_{t}^{k}}s_{\phi}(\mathbf{O}_{t},\mathbf{A}_{t}^{k})$ (12)
@@ V-A2 Tasks
As shown in Fig. 5, we design five challenging contact-rich tasks. To thoroughly evaluate the performance of our paradigm, these tasks include long-horizon, precise manipulation (e.g., Shoe Lacing), collaborative handover of fragile objects (e.g., Chip Handover), and low-data (only 20 demos), high-precision manipulation (e.g., Lock Opening). More detailed task descriptions and examples of the evaluation metrics are provided in Appendix -H.
@@ V-A3 Baselines
We adopt two state-of-the-art (SOTA) visuomotor policies (i.e., DP [14] and $\pi_{0.5}$ [32]) as the base policy for TouchGuide, and use several visuo-tactile policies (e.g., RDP [65], Policy Consensus [9] and SafeDiff [57]) as the baselines for comparison. Specifically, we include Tactile Dynamics as one of our baselines. This method is derived from Latent Policy Barrier [51] and employs a dynamics model to jointly predict tactile and visual observations. Details on the touch guidance setup of Tactile Dynamics for $\pi_{0.5}$ can be found in Sec. IV-A. For TouchGuide, we use two tactile modalities (i.e., force and tactile image) to thoroughly validate the cross-modality capability of our paradigm. The specific baseline configurations are detailed in Appendix -G. For these five tasks, each baseline is evaluated over 20 trials with moderate randomization of the initial environment conditions, and the results are reported in Table I.

### v2

@@ Abstract
Fine-grained and contact-rich manipulation remain challenging for robots, largely due to the underutilization of tactile feedback. To address this, we introduce TouchGuide, a novel cross-policy visuo-tactile fusion paradigm that fuses modalities within a low-dimensional action space. Specifically, TouchGuide operates in two stages to guide a pre-trained diffusion or flow-matching visuomotor policy at inference time. First, the policy produces a coarse, visually-plausible action using only visual inputs during early sampling. Second, a task-specific Contact Physical Model (CPM) provides touch guidance to steer and refine the action, ensuring it aligns with realistic physical contact conditions. Trained through contrastive learning on limited expert demonstrations, the CPM provides a tactile-informed feasibility score to steer the sampling process toward refined actions that satisfy physical contact constraints. Furthermore, to facilitate TouchGuide training with high-quality and cost-effective data, we introduce TacUMI, a data collection system. TacUMI achieves a favorable trade-off between precision and affordability; by leveraging rigid fingertips, it obtains direct tactile feedback, thereby enabling the collection of reliable tactile data. Extensive experiments on five challenging contact-rich tasks, such as shoe lacing and chip handover, show that TouchGuide consistently and significantly outperforms state-of-the-art visuo-tactile policies.
@@ I Introduction
To address this fundamental limitation, we introduce TouchGuide, a novel cross-policy visuo-tactile fusion paradigm. Departing from conventional computer vision approaches that focus on “fusing multi-modal data per se”, we rethink the problem from a robotic control perspective. Our core idea is to extract task-relevant semantics from each modality and perform fusion within the low-dimensional action space in a way that can benefits robotic manipulation. Concretely, TouchGuide steers a pre-trained visuomotor policy based on diffusion or flow matching (e.g., the action expert in $\pi_{0.5}$) during inference. The process begins by generating a visually feasible, coarse action from visual observation alone in the early sampling phase. Then, in the later phase, a task-specific Contact Physical Model (CPM) injects touch guidance, refining the action to better satisfy real-world physical contact constraints. The CPM is trained via contrastive learning on limited expert demonstrations to directly output a feasibility score. This score assesses whether the action comply with the contact physical constraints implied by the current tactile observations in task-specific environment, thereby generating a more robust steering signal. Notably, CPM can be applied across multiple policies, including VLAs [32, 5, 50, 58], diffusion policies [14, 69, 56, 44], and flow-matching policies [34, 71].
Moreover, in fine-grained and contact-rich manipulation, policy execution performance depends heavily on the quality of expert demonstrations. Teleoperation, due to high latency and indirect haptic feedback, hinders intuitive control. While handheld devices such as UMI [15] provide a platform-decoupled solution for data collection, existing systems relying on SLAM-based localization or motion capture (MoCap) cannot simultaneously achieve both high accuracy and low cost. To this end, we introduce TacUMI, designed according to the following principles: (1) Cost-effective (base cost \$720, tactile sensors are additional, using Xense sensors) and lightweight (540 g), while providing the high-precision data required for fine-grained manipulation. (2) Direct tactile feedback enabled by rigid fingertips. (3) A unified interface for synchronized acquisition of high-resolution visual and tactile signals.
@@ II Related Work
Figure 2: Overview of TacUMI data collection system. (a) TacUMI (Collection-side) uses a Vive tracker for localization to obtain accurate end-effector poses, while the operator receives direct tactile feedback. (b) During policy inference, we use an execution-side device that is structurally identical to the collection-side TacUMI, coupled to different robot arms via an adapter.
@@ III-A System Overview
Accuracy and Low-cost. TacUMI is localized using a Vive tracker with two Lighthouse base stations, providing reliable and accurate pose estimates. In contrast, UMI [15] relies on SLAM-based tracking; for fine-grained tactile manipulation tasks which requires high precision, its accuracy is insufficient and tracking can be easily lost in feature-sparse scenes. Notably, a setup with one Vive tracker and two Lighthouse base stations has a base cost of only about \$720, maintaining low cost while providing accurate localization. A detailed comparison of data collection quality across different methods is provided in the experiments (see Sec. V-D).
@@
Usability and Synchronization. TacUMI adopts a unified hardware-software interface. With a single aviation cable, it streams all observations in real time, including visual and tactile signals and proprioceptive states. Our data collection is compatible with the LeRobot [7] format and uses one software pipeline to acquire all modalities in a unified timestamped stream, enabling effective software-level synchronization. TacUMI captures visual $\mathbf{V}_{t}\in\mathbb{R}^{640\times 480\times 3}$ and tactile images $\mathbf{T}_{t}^{\text{image}}\in\mathbb{R}^{200\times 350\times 3}$ at 30Hz, sufficient for most fine-grained manipulation tasks. We also provide force observations $\mathbf{T}_{t}^{\text{force}}\in\mathbb{R}^{20\times 35\times 3}$ as a dense 3D force field; in our algorithm, it is handled the same as tactile images, except for the input dimensionality. An onboard RK3576 CPU runs the XenseSDK [61] to compute this force field online from the tactile images, without requiring external processing, as data are being captured. Since tactile sensing is modality-specific (i.e., tactile images are not shared across sensors for training), this capability is crucial for scalable tactile data collection.
More details on hardware selection, design choices, and system parameters are provided in the Appendix -E.
@@ III-B Data Processing for Training
$\mathbf{p}_{\text{rel}}^{(i)}$ $={\mathbf{R}_{\text{ee}}^{(0)\top}}(\mathbf{p}_{\text{ee}}^{(i)}-\mathbf{p}_{\text{ee}}^{(0)}),$ (1)
$\mathbf{R}_{\text{rel}}^{(i)}$ $={\mathbf{R}_{\text{ee}}^{(0)\top}}\mathbf{R}_{\text{ee}}^{(i)},$ (2)
@@ IV Method: TouchGuide
In this section, we introduce TouchGuide, a two-stage cross-policy multimodal fusion paradigm based on classifier guidance [18] (Sec. IV-A), inspired by recent advances in inference-time steering [51, 22, 55]. TouchGuide performs inference-time steering on a pre-trained diffusion-based or flow-matching-based (e.g., action expert of $\pi_{0.5}$) visuomotor policy $\pi_{\theta}(\mathbf{A}_{t}\,|\,\mathbf{V}_{t})$. During the early phase of the denoising or flow-matching process, the policy generates an initial coarse action based solely on visual observation, ensuring visual feasibility without yet enforcing tactile constraints. In the later phase, a task-specific Contact Physical Model (CPM, detailed in Sec. IV-B) provides touch guidance via a feasibility score, which steers the action toward better satisfaction of real-world physical contact. The complete inference-time guidance process of TouchGuide is detailed in Sec. IV-C.
@@ IV-A Preliminaries: Classifier Guidance
$\hat{\epsilon}_{\theta}(x_{t})=\epsilon_{\theta}(x_{t})-\eta\sqrt{1-\bar{\alpha}_{t}}\nabla_{x_{t}}\log p_{\phi}(y\mid x_{t}).$ (3)
@@
$\hat{u}_{\theta}(x_{t})=u_{\theta}(x_{t})-\eta\frac{t}{1-t}\nabla_{x_{t}}\log p_{\phi}(y\mid x_{t}),$ (4)
@@ IV-B Contact Physical Model
At time step $t$, Contact Physical Model (CPM, see Fig. 3(a)) outputs a task-specific feasibility score conditioned on the tactile feedback $\mathbf{T}_{t}$, the visual observation $\mathbf{V}_{t}$ and the coarse action $\mathbf{A}_{t}$ generated by the early phase of pre-trained policy $\pi_{\theta}(\mathbf{A}_{t}\,|\,\mathbf{V}_{t})$. Specifically, we use pretrained DINOv2 as both the tactile and visual encoder to obtain the tactile and visual embedings, i.e., $\mathbf{T}_{t}^{\text{emb}}$ and $\mathbf{V}^{\text{emb}}_{t}$ respectively, as it provides effective representations of robotic observations [22, 78] (see Appendix -C for encoder comparison experiments). Then, we fuse these two embedings using N-layer Transformer encoder $\mathtt{TranEnc}(\cdot)_{\times\text{N}}$ and finally, yield the latent observation $\mathbf{O}_{t}$:
$\mathbf{O}_{t}^{\text{emb}}=\mathtt{TranEnc}(\mathtt{concat}[\mathbf{T}_{t}^{\text{emb}},\mathbf{V}_{t}^{\text{emb}}])_{\times\text{N}}\xrightarrow[]{\text{L2 Norm}}\mathbf{O}_{t}.$ (5)
Similarly, we can obtain the latent action $\mathbf{a}_{t}$ by:
$\mathbf{A}_{t}^{\text{emb}}=\mathcal{E}_{\mathbf{A}}(\mathbf{A}_{t})\xrightarrow[]{\text{L2 Norm}}\mathbf{a}_{t},$ (6)
where $\mathcal{E}_{\mathbf{A}}$ is the action encoder (1D CNN + MLP) and $\mathbf{A}_{t}^{\text{emb}}$ is the action embedding. The feasibility score $\mathbf{s}$ represents the cosine similarity between the observation and the action in the latent space [11], and is defined as follows:
$\mathbf{s}=s_{\phi}(\mathbf{V}_{t},\mathbf{T}_{t},\mathbf{A}_{t})=\mathbf{O}_{t}^{\top}\mathbf{a}_{t}.$ (7)
$\mathcal{L}_{\text{CPM}}(\mathbf{O}_{t},\mathbf{a}_{t})=\frac{1}{2}(\mathcal{L}_{\mathbf{O}\to\mathbf{a}}+\mathcal{L}_{\mathbf{a}\to\mathbf{O}}),$ (8)
$\mathcal{L}_{\mathbf{O}\to\mathbf{a}}$ $=-\frac{1}{M}\sum_{i=1}^{M}\log\frac{\exp(\mathbf{O}_{i}^{\top}\mathbf{a}_{i}/\tau)}{\sum_{j=1}^{M}\exp(\mathbf{O}_{i}^{\top}\mathbf{a}_{j}/\tau)},$ (9)
$\mathcal{L}_{\mathbf{a}\to\mathbf{O}}$ $=-\frac{1}{M}\sum_{i=1}^{M}\log\frac{\exp(\mathbf{a}_{i}^{\top}\mathbf{O}_{i}/\tau)}{\sum_{j=1}^{M}\exp(\mathbf{a}_{i}^{\top}\mathbf{O}_{j}/\tau)},$ (10)
where $\tau$ is a learnable temperature parameter and $M$ denotes the number of samples used during training. In order to enable the CPM to effectively operate in a noisy action space, which is generated by the early phase of the pretrained policy during inference, we enhance the ground-truth action $\mathbf{A}_{t}$ by adding noise following a geometric distribution using the same noise scheduler as DP [22]; for flow-matching policies, we instead apply linear interpolation.
@@
18:     $\hat{u}\leftarrow u-\eta\frac{k}{1-k}\nabla_{\mathbf{A}_{t}^{k}}\mathbf{s}$
@@ IV-C Steering the Policies via Touch Guidance
$\hat{\epsilon}_{\theta}=\epsilon_{\theta}(\mathbf{A}_{t}^{k},\mathbf{V}_{t})-\eta\sqrt{1-\bar{\alpha}_{k}}\nabla_{\mathbf{A}_{t}^{k}}s_{\phi}(\mathbf{V}_{t},\mathbf{T}_{t},\mathbf{A}_{t}^{k}),$ (11)
$\hat{u}_{\theta}=u_{\theta}(\mathbf{A}_{t}^{k},\mathbf{V}_{t})-\eta\frac{k}{1-k}\nabla_{\mathbf{A}_{t}^{k}}s_{\phi}(\mathbf{O}_{t},\mathbf{T}_{t}\mathbf{A}_{t}^{k}),$ (12)
@@ V-A2 Tasks
As shown in Fig. 5, we design five challenging contact-rich tasks. To thoroughly evaluate the performance of our paradigm, these tasks include long-horizon, precise manipulation (e.g., Shoe Lacing), collaborative handover of fragile objects (e.g., Chip Handover), and low-data (only 20 demos), high-precision manipulation (e.g., Lock Opening). More detailed task descriptions and examples of the evaluation metrics are provided in Appendix -I.
@@ V-A3 Baselines
We adopt two state-of-the-art (SOTA) visuomotor policies (i.e., DP [14] and $\pi_{0.5}$ [32]) as the base policy for TouchGuide, and use several visuo-tactile policies (e.g., RDP [65], Policy Consensus [9] and SafeDiff [57]) as the baselines for comparison. Specifically, we include Tactile Dynamics as one of our baselines. This method is derived from Latent Policy Barrier [51] and employs a dynamics model to jointly predict tactile and visual observations. Details on the touch guidance setup of Tactile Dynamics for $\pi_{0.5}$ can be found in Sec. IV-A. For TouchGuide, we use two tactile modalities (i.e., force and tactile image) to thoroughly validate the cross-modality capability of our paradigm. The specific baseline configurations are detailed in Appendix -H. For these five tasks, each baseline is evaluated over 20 trials with moderate randomization of the initial environment conditions, and the results are reported in Table I.